OpenStudy (anonymous):

Solve the following system x-3y = 10 2x+y = 13

OpenStudy (anonymous):

x-3y = 10--> x = 10+3y Plug in for x.....2(10+3y) + y = 13---->20 + 7y = 13 Solving for y gives -1. Plug that in for y and solve it.

OpenStudy (anonymous):

x-3y = 10 2x+y = 13 2{x-3y}=(2)(10) -(2x+y)=(-1)(13) 2x - 6y = 20 -2x-y = -13 add them up -7y = 7 y = -1 substitute y =-1 to one of them x-3(-1)=10 x+4 = 10 x = 6

OpenStudy (anonymous):

So the answer is x=7 y=-1 Correct?

OpenStudy (anonymous):

@franciscanmonk I think you made a mistake here x-3(-1)=10

OpenStudy (anonymous):

Yep, that's correct joker @thejoker1177